Find an equation of a parabola with focus(5,0) and directrix x = -5

Thanks so much in advance:)

Answer by TimothyLamb(4379)

---
given:
focus(5,0)
directrix x = -5
---
directrix is a vertical line, so parabola is horizontal.
---
find vertex, half way between focus and directrix:
---
h = (5 - (-5))/2 + -5
h = (5 + 5)/2 + -5
h = 5 + -5
h = 0
vertex = (h,k) = (0,0)
---
directrix is to the left of the focus so parabola opens to the right and p is positive
distance between focus and vertex: p = 5
---
conic form of horizontal parabola:
4p(x – h) = (y – k)^2
4(5)x = y^2
---
answer:
20x = y^2
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php